OpenStudy (anonymous):

Let f(x) = 4x – 32 and g(x) = 4x. Part 1 What is (f ◦ g)(x) ? Part 2 Use complete sentences to describe the method you used to solve this problem.

OpenStudy (anonymous):

(f ◦ g)(x) = f(g(x)) = f(4x) = 4(4x) - 32 = 16x - 32

OpenStudy (anonymous):

ok?

OpenStudy (anonymous):

okk

OpenStudy (anonymous):

Part 2?

OpenStudy (anonymous):

how would you explain it?

OpenStudy (anonymous):

put the function g into f and simplify... this is called composition of functions.

OpenStudy (anonymous):

oh okay, i get it. thank you